OpenStudy (deadshot):

how do I solve logv4 (y - 9) + logv4 3 = logv4 81?

OpenStudy (deadshot):

\[\log _{4} (y-9)+\log _{4} 3=\log _{4} 81\]

OpenStudy (kropot72):

First you use the following rule of logs on the left hand side: \[\log_{a} b+\log_{a} c=\log_{a} (b \times c)\] Can you do that?

OpenStudy (deadshot):

so \[\log _{4} (3y-27)=\log _{4} 81\]

OpenStudy (kropot72):

Good work! Now we can state the following: 3y - 27 = 81 and solve for y.

OpenStudy (kropot72):

@DeadShot Are you there?

OpenStudy (deadshot):

Sorry, I'm taking care of my sick little brother while my parents are away, so y=36, right?

OpenStudy (kropot72):

Your answer is correct :)

OpenStudy (deadshot):

okay, thanks!

OpenStudy (kropot72):

You're welcome :)